# Statistical-Computational Trade-offs for Recursive Adaptive Partitioning Estimators

### v2

@@ Appendix F Stable MSP Properties
$\Delta(k_{0}(\mathcal{C});\mathcal{C},\mathcal{D}_{n})\geq 2^{\left|J(\mathcal{C})\right|}\lambda$ (G.20)
for some choice of $k_{0}(\mathcal{C})\in S^{*}$. Using Lemma G.1 together with (47) and a union bound, we therefore get a probability at least $31-\delta/$ 3 event on which
$\begin{split}\hat{\Delta}^{1/2}(k;\mathcal{C},\mathcal{D}_{n})&\geq\Delta^{1/2}(k;\mathcal{C},\mathcal{D}_{n})-\sqrt{\frac{18(9\|f^{*}\|_{\infty}^{2}+\sigma^{2})\left((s+2)\log 3+\log(2/\delta)\right)}{N(\mathcal{C})}}\\ &\geq\sqrt{2^{|J(\mathcal{C})|}\lambda}-\sqrt{\frac{18(9\|f^{*}\|_{\infty}^{2}+\sigma^{2})\left((s+2)\log 3+\log(2/\delta)\right)}{N(\mathcal{C})}}\end{split}$ (G.21)
uniformly over all $\mathcal{C}\in{\mathfrak{C}^{*}}$. Reasoning similarly, we get another probability at least $31-\delta/$ 3 event on which
$\max_{k\notin S^{*}}\hat{\Delta}^{1/2}(k;\mathcal{C},\mathcal{D}_{n})\leq\sqrt{\frac{18(9\|f^{*}\|_{\infty}^{2}+\sigma^{2})\left((s+2)\log 3+\log(2d/\delta)\right)}{N(\mathcal{C})}}$ (G.22)
uniformly over all $\mathcal{C}\in\mathfrak{C}$. Finally, using (G.9) together with a union bound, we get (G.15). ∎
@@
Theorem G.3.
Under Assumption 3.1, let $f_{0}^{*}\in\textsf{SMSP}(\lambda)$. Suppose $\varepsilon$ is sub-Gaussian with parameter $\sigma$ (Wainwright, 2019). Fix $10<\delta<$ 1. Denote
$\tau\coloneqq 18(9\|f^{*}\|_{\infty}^{2}+\sigma^{2})\left((s+2)\log 3+\log(2d/\delta)\right)$ (G.23)
@@
$R\left(\hat{f}_{\operatorname{CART}}(-;\mathcal{D}_{n}),f^{*}\right)\leq\frac{2^{s+1}(\|f^{*}\|_{\infty}^{2}+\sigma^{2})\left((s+1)\log 3+\log(2/\delta)\right)}{n}.$ (G.24)
@@
$\mathfrak{R}\left(\hat{f}_{\operatorname{CART}},f^{*}_{0},d,n\right)=O\left(\frac{2^{s}(\|f^{*}\|_{\infty}^{2}+\sigma^{2})(s+\log n)}{n}\right).$ (G.25)
@@
We will prove (G.24) in three steps. The first two steps show that with high probability, the tree $\mathfrak{T}_{\operatorname{CART}}$ fitted using CART with the stopping condition specified by Theorem 7.5 does not make any splits on irrelevant covariates and that $f^{*}$ is constant on its leaves. The third step will show that the mean responses over the leaves of $\mathfrak{T}_{\operatorname{CART}}$ concentrate around their true values with high probability. These together will give us the result that we want.
Step 1: No bad splits. Let $\Omega$ denote the event on which (G.14), (G.13) and (G.15) hold. We first claim that on $\Omega$, $\mathfrak{T}_{\operatorname{CART}}$ does not make any splits in irrelevant covariates. Suppose not, then there exists a split on an irrelevant covariate occurring at a minimum depth. By minimality, this split must occur on a cell $\mathcal{C}$ in $\mathfrak{C}$. On $\Omega$, we have by equation (G.14) that
$\max_{k\notin S^{*}}\frac{N(\mathcal{C})}{n}\hat{\Delta}(k;\mathcal{C},\mathcal{D}_{n})\leq\frac{18(9\|f^{*}\|_{\infty}^{2}+\sigma^{2})\left((s+2)\log 3+\log(2d/\delta)\right)}{n}<\gamma,$ (G.26)
@@
Step 2: $f^{*}$ is constant on leaves. We have already shown that every leaf $\mathcal{C}$ of $\mathfrak{T}_{\operatorname{CART}}$ is an element of $\mathfrak{C}$. We now claim that, upon further conditioning on the event on which (G.16) holds, $\mathcal{C}$ must also be an element of $\superscript\mathfrak{C}\backslash{\mathfrak{C}^{*}}$. Suppose not and instead $\mathcal{C}\in{\mathfrak{C}^{*}}$, then by combining equations (G.13) and (G.16), we get
$\begin{split}\max_{k\in S^{*}}\frac{N(\mathcal{C})}{n}\hat{\Delta}(k;\mathcal{C},\mathcal{D}_{n})&\geq\left(\sqrt{\frac{N(\mathcal{C})2^{|J(\mathcal{C})|}\lambda}{n}}-\sqrt{\frac{18(9\|f^{*}\|_{\infty}^{2}+\sigma^{2})\left((s+2)\log 3+\log(2d/\delta)\right)}{n}}\right)_{+}^{2}\\ &\geq\left(\sqrt{\frac{\lambda}{2}}-\sqrt{\frac{18(9\|f^{*}\|_{\infty}^{2}+\sigma^{2})\left((s+2)\log 3+\log(2d/\delta)\right)}{n}}\right)_{+}^{2}\\ &>\gamma,\end{split}$ (G.27)
@@
Step 3: Concentration of function values. Let $\mathfrak{L}$ denote the set of leaves of $\mathfrak{T}_{\operatorname{CART}}$. By the previous two steps, we know that $\𝔏\superscript\mathfrak{L}\subset\mathfrak{C}\backslash{\mathfrak{C}^{*}}$, which implies that $\left|J(\mathcal{C})\right|\leq s$ and that $f^{*}(\mathbf{x})=\mathbb{E}\left\{Y~{}|~{}\mathbf{x}\in\mathcal{C}\right\}$ for all $\mathcal{C}\in\mathfrak{L}$. As a consequence of (G.15), for any $\mathbf{x}\in\{\pm 1\}^{d}$ and letting $\mathcal{C}(\mathbf{x})$ be the leaf that contains it, we have
$\begin{split}\left|\hat{f}_{\operatorname{CART}}(\mathbf{x})-f^{*}(\mathbf{x})\right|&=\left|\bar{Y}_{I(\mathcal{C})}-f^{*}(\mathbf{x})\right|\\ &=\left|\bar{Y}_{I(\mathcal{C})}-\mathbb{E}\left\{Y~{}|~{}\mathbf{X}\in\mathcal{C}(\mathbf{x})\right\}\right|\\ &\leq\sqrt{\frac{2(\|f^{*}\|_{\infty}^{2}+\sigma^{2})\left((s+1)\log 3+\log(2/\delta)\right)}{N(\mathcal{C}(\mathbf{x}))}}\\ &\leq\sqrt{\frac{2^{s+1}(\|f^{*}\|_{\infty}^{2}+\sigma^{2})\left((s+1)\log 3+\log(2/\delta)\right)}{n}}.\end{split}$ (G.28)
@@
$R\left(\hat{f}_{\operatorname{CART}}(-;\mathcal{D}_{n}),f^{*}\right)\leq\frac{2^{s+1}(\|f^{*}\|_{\infty}^{2}+\sigma^{2})\left((s+1)\log 3+\log(2/\delta)\right)}{n},$ (G.29)
@@
To prove (G.25), we take $\delta=1/n$ in (G.24). Let $\mathcal{E}$ be the event on which (G.24) holds. We then write
$\mathfrak{R}\left(\hat{f}_{\operatorname{CART}},f^{*}_{0},d,n\right)=\mathbb{E}\left\{R\left(\hat{f}_{\operatorname{CART}}(-;\mathcal{D}_{n}),f^{*}\right)\mathbb{1}_{\mathcal{E}}\right\}+\mathbb{E}\left\{R\left(\hat{f}_{\operatorname{CART}}(-;\mathcal{D}_{n}),f^{*}\right)\mathbb{1}_{\mathcal{E}^{c}}\right\}.$ (G.30)
@@
$\|\hat{f}_{\operatorname{CART}}(-;\mathcal{D}_{n})\|_{\infty}\leq\max_{1\leq i\leq n}\left|Y_{i}\right|,$ (G.31)
@@
$R\left(\hat{f}_{\operatorname{CART}}(-;\mathcal{D}_{n}),f^{*}\right)\leq\left(\max_{1\leq i\leq n}\left|Y_{i}\right|+\left\|f^{*}\right\|_{\infty}\right)^{2}.$ (G.32)
@@
$\max_{1\leq i\leq n}\left|Y_{i}\right|\lesssim\sigma\left(\sqrt{\log n}+t\right)$ (G.33)
with probability at least $1-2e^{-t^{2}}$, for any $t>0$. Plugging this into (G.32) and applying Cauchy-Schwarz twice gives
$R\left(\hat{f}_{\operatorname{CART}}(-;\mathcal{D}_{n}),f^{*}\right)\lesssim\left\|f^{*}\right\|_{\infty}^{2}+\sigma^{2}\log n+\sigma^{2}u$ (G.34)
@@
$\mathbb{P}\left\{R\left(\hat{f}_{\operatorname{CART}}(-;\mathcal{D}_{n}),f^{*}\right)\geq u\right\}\leq 2e^{-C(u-\|f^{*}\|_{\infty}^{2}+\sigma^{2}\log n)/\sigma^{2}}$ (G.35)
@@
[…]
This completes the bound on the second term and hence the proof of (G.25). ∎
Appendix H Proof of Theorem 8.1
@@
Imitating the proof of Theorem 5.1, we combine Lemma C.2 together with an extension of Lemma C.1 to accommodate a vertex cut and the new coupling argument. ∎
Lemma H.1.
@@
$\operatorname{KL}\left(\mathcal{D}_{n}\|\check{\mathcal{D}}_{n}\right)=\sum_{i=1}^{n}\operatorname{KL}\left((\mathbf{X}_{i},Y_{i})\|(\mathbf{X}_{i},\check{Y}_{i})\right).$ (H.1)
@@
$\operatorname{KL}\left((\mathbf{X}_{i},Y_{i})\|(\mathbf{X}_{i},\check{Y}_{i})\right)=\operatorname{KL}(\mathbf{X}_{i}\|\mathbf{X}_{i})+\mathbb{E}\left\{\operatorname{KL}\left((Y_{i}|\mathbf{X}_{i})\|(\check{Y}_{i}|\mathbf{X}_{i})\right)\right\}.$ (H.2)
The first term on the right hand side is equal to 00 since the covariate distributions are the same. To compute the second term, notice that $Y_{i}|\mathbf{X}_{i}\sim\mathcal{N}(f^{*}(\mathbf{X}_{i}),\sigma^{2})$, and $\check{Y}_{i}|\mathbf{X}_{i}\sim\mathcal{N}(\check{f}^{*}(\mathbf{X}_{i}),\sigma^{2})$. The KL divergence between them is thus
$\operatorname{KL}\left((Y_{i}|\mathbf{X}_{i})\|(\check{Y}_{i}|\mathbf{X}_{i})\right)=\frac{\left(f^{*}(\mathbf{X}_{i})-\check{f}^{*}(\mathbf{X}_{i})\right)^{2}}{2\sigma^{2}}.$ (H.3)
@@
$\mathbb{E}\left\{\operatorname{KL}\left((Y_{i}|\mathbf{X}_{i})\|(\check{Y}_{i}|\mathbf{X}_{i})\right)\right\}=\frac{\mathbb{E}\left\{\left(\sum_{S\in\mathcal{B}}\alpha_{S}\chi_{S}(\mathbf{X}_{i})\right)\right\}^{2}}{2\sigma^{2}}=\frac{w(\mathcal{B})}{2\sigma^{2}}.$ (H.4)
Plugging this back into (H.1) gives
$\operatorname{KL}\left(\mathcal{D}_{n}\|\check{\mathcal{D}}_{n}\right)\leq\frac{nw(\mathcal{B})}{2\sigma^{2}},$ (H.5)
@@
$\operatorname{TV}\left(\mathcal{D}_{n},\check{\mathcal{D}}_{n}\right)\leq\sqrt{\frac{nw(\mathcal{B})}{4\sigma^{2}}}.$ (H.6)
@@
$\begin{split}\mathbb{P}_{\mathcal{D}_{n}}\left\{J(\mathbf{x};\mathcal{D}_{n},\Theta)\cap T\neq\emptyset\right\}&\leq\mathbb{P}_{\mathcal{D}_{n}}\left\{J(\mathbf{x};\check{\mathcal{D}}_{n},\Theta)\cap T\neq\emptyset~{}\text{or}~{}\check{\mathcal{D}}_{n}\neq\mathcal{D}_{n}\right\}\\ &\leq\mathbb{P}_{\mathcal{D}_{n}}\left\{J(\mathbf{x};\check{\mathcal{D}}_{n},\Theta)\cap T\neq\emptyset\right\}+\mathbb{P}\left\{\check{\mathcal{D}}_{n}\neq\mathcal{D}_{n}\right\}\\ &\frac{s_{\operatorname{T}}(\log_{2}n+2)}{d-s_{-\mathcal{B},\textsf{MSP}}-s_{\operatorname{T}}+1}+\sqrt{\frac{nw(\mathcal{B})}{4\sigma^{2}}}.\qed\end{split}$ (H.7)
Appendix I Constant Probability Lower Bounds
@@
$R\left(\hat{f}_{\operatorname{G}}(-;\mathcal{D}_{n},\Theta),f^{*}\right)\geq\frac{\operatorname{Var}\left\{r_{\textsf{MSP}}(\mathbf{X})\right\}}{6}$ (I.1)
@@
$\log_{2}n\leq\frac{\operatorname{Var}\{r_{\textsf{MSP}}(\mathbf{X})\}(d-s_{\textsf{MSP}}-s_{\operatorname{T}}+1)}{8M^{2}s_{\textsf{MSP}}}-2,$ (I.2)
@@
$\mathbb{E}_{\Theta}\left\{R\left(\hat{f}_{\operatorname{G-E}}(-;\mathcal{D}_{n},\Theta)\right)\right\}\geq\frac{\operatorname{Var}\left\{r_{\textsf{MSP}}(\mathbf{x})\right\}}{4}.$ (I.3)
@@
$R\left(\hat{f}_{\operatorname{G}}(-;\mathcal{D}_{n},\Theta),f^{*}\right)\geq V(\mathcal{D}_{n},\Theta)\operatorname{Var}\left\{r_{\textsf{MSP}}(\mathbf{X})\right\},$ (I.4)
@@
$V(\mathcal{D}_{n},\Theta)\coloneqq\frac{\mathbb{E}_{\mathbf{X}_{[d]\backslash T}}\left\{\operatorname{Var}\left\{f^{*}(\mathbf{X})~{}|~{}\mathbf{X}_{[d]\backslash T}\right\}\mathbb{1}\left\{J(\mathbf{X};\mathcal{D}_{n},\Theta)\cap T=\emptyset\right\}\right\}}{\mathbb{E}_{\mathbf{X}_{[d]\backslash T}}\left\{\operatorname{Var}\left\{f^{*}(\mathbf{X})~{}|~{}\mathbf{X}_{[d]\backslash T}\right\}\right\}}.$ (I.5)
It is clear that $0\leq V(\mathcal{D}_{n},\Theta)\leq 1$. Furthermore, using Lemma C.1, we see that $\mathbb{E}_{\mathcal{D}_{n}}\left\{V(\mathcal{D}_{n},\Theta)\right\}\geq 1-\delta$, where $\delta\coloneqq\frac{s_{\textsf{MSP}}(\log_{2}n+2)}{d-s_{\textsf{MSP}}-s_{\operatorname{T}}+1}$. Next, making use of the simple inequality
$\mathbb{E}_{\mathcal{D}_{n}}\left\{V(\mathcal{D}_{n},\Theta)\right\}\leq\mathbb{P}_{\mathcal{D}_{n}}\left\{V(\mathcal{D}_{n},\Theta)\geq\delta\right\}+\delta,$ (I.6)
@@
$\mathbb{P}_{\mathcal{D}_{n}}\left\{V(\mathcal{D}_{n})\geq\delta\right\}\geq 1-3\delta.$ (I.7)
@@
$R\left(\hat{f}_{\operatorname{G}}(-;\mathcal{D}_{n},\Theta),f^{*}\right)\geq\frac{\operatorname{Var}\left\{r_{\textsf{MSP}}(\mathbf{X})\right\}}{6}.$ (I.8)

### v3

@@ Appendix G Stable MSP Properties
$\Delta(k_{0}(\mathcal{C});\mathcal{C},\mathcal{D}_{n})\geq 2^{\left|J(\mathcal{C})\right|}\lambda$ (H.20)
for some choice of $k_{0}(\mathcal{C})\in S^{*}$. Using Lemma H.1 together with (46) and a union bound, we therefore get a probability at least $1-\delta/3$ event on which
$\begin{split}\hat{\Delta}^{1/2}(k;\mathcal{C},\mathcal{D}_{n})&\geq\Delta^{1/2}(k;\mathcal{C},\mathcal{D}_{n})-\sqrt{\frac{18(9\|f^{*}\|_{\infty}^{2}+\sigma^{2})\left((s+2)\log 3+\log(2/\delta)\right)}{N(\mathcal{C})}}\\ &\geq\sqrt{2^{|J(\mathcal{C})|}\lambda}-\sqrt{\frac{18(9\|f^{*}\|_{\infty}^{2}+\sigma^{2})\left((s+2)\log 3+\log(2/\delta)\right)}{N(\mathcal{C})}}\end{split}$ (H.21)
uniformly over all $\mathcal{C}\in{\mathfrak{C}^{*}}$. Reasoning similarly, we get another probability at least $1-\delta/3$ event on which
$\max_{k\notin S^{*}}\hat{\Delta}^{1/2}(k;\mathcal{C},\mathcal{D}_{n})\leq\sqrt{\frac{18(9\|f^{*}\|_{\infty}^{2}+\sigma^{2})\left((s+2)\log 3+\log(2d/\delta)\right)}{N(\mathcal{C})}}$ (H.22)
uniformly over all $\mathcal{C}\in\mathfrak{C}$. Finally, using (H.9) together with a union bound, we get (H.15). ∎
@@
Theorem H.3.
Under Assumption 3.1, let $f_{0}^{*}\in\textsf{SMSP}(\lambda)$. Suppose $\varepsilon$ is sub-Gaussian with parameter $\sigma$ (Wainwright, 2019). Fix $0<\delta<1$. Denote
$\tau\coloneqq 18(9\|f^{*}\|_{\infty}^{2}+\sigma^{2})\left((s+2)\log 3+\log(2d/\delta)\right)$ (H.23)
@@
$R\left(\hat{f}_{\operatorname{CART}}(-;\mathcal{D}_{n}),f^{*}\right)\leq\frac{2^{s+1}(\|f^{*}\|_{\infty}^{2}+\sigma^{2})\left((s+1)\log 3+\log(2/\delta)\right)}{n}.$ (H.24)
@@
$\mathfrak{R}\left(\hat{f}_{\operatorname{CART}},f^{*}_{0},d,n\right)=O\left(\frac{2^{s}(\|f^{*}\|_{\infty}^{2}+\sigma^{2})(s+\log n)}{n}\right).$ (H.25)
@@
We will prove (H.24) in three steps. The first two steps show that with high probability, the tree $\mathfrak{T}_{\operatorname{CART}}$ fitted using CART with the stopping condition specified by Theorem 7.5 does not make any splits on irrelevant covariates and that $f^{*}$ is constant on its leaves. The third step will show that the mean responses over the leaves of $\mathfrak{T}_{\operatorname{CART}}$ concentrate around their true values with high probability. These together will give us the result that we want.
Step 1: No bad splits. Let $\Omega$ denote the event on which (H.14), (H.13) and (H.15) hold. We first claim that on $\Omega$, $\mathfrak{T}_{\operatorname{CART}}$ does not make any splits in irrelevant covariates. Suppose not, then there exists a split on an irrelevant covariate occurring at a minimum depth. By minimality, this split must occur on a cell $\mathcal{C}$ in $\mathfrak{C}$. On $\Omega$, we have by equation (H.14) that
$\max_{k\notin S^{*}}\frac{N(\mathcal{C})}{n}\hat{\Delta}(k;\mathcal{C},\mathcal{D}_{n})\leq\frac{18(9\|f^{*}\|_{\infty}^{2}+\sigma^{2})\left((s+2)\log 3+\log(2d/\delta)\right)}{n}<\gamma,$ (H.26)
@@
Step 2: $f^{*}$ is constant on leaves. We have already shown that every leaf $\mathcal{C}$ of $\mathfrak{T}_{\operatorname{CART}}$ is an element of $\mathfrak{C}$. We now claim that, upon further conditioning on the event on which (H.16) holds, $\mathcal{C}$ must also be an element of $\mathfrak{C}\backslash{\mathfrak{C}^{*}}$. Suppose not and instead $\mathcal{C}\in{\mathfrak{C}^{*}}$, then by combining equations (H.13) and (H.16), we get
$\begin{split}\max_{k\in S^{*}}\frac{N(\mathcal{C})}{n}\hat{\Delta}(k;\mathcal{C},\mathcal{D}_{n})&\geq\left(\sqrt{\frac{N(\mathcal{C})2^{|J(\mathcal{C})|}\lambda}{n}}-\sqrt{\frac{18(9\|f^{*}\|_{\infty}^{2}+\sigma^{2})\left((s+2)\log 3+\log(2d/\delta)\right)}{n}}\right)_{+}^{2}\\ &\geq\left(\sqrt{\frac{\lambda}{2}}-\sqrt{\frac{18(9\|f^{*}\|_{\infty}^{2}+\sigma^{2})\left((s+2)\log 3+\log(2d/\delta)\right)}{n}}\right)_{+}^{2}\\ &>\gamma,\end{split}$ (H.27)
@@
Step 3: Concentration of function values. Let $\mathfrak{L}$ denote the set of leaves of $\mathfrak{T}_{\operatorname{CART}}$. By the previous two steps, we know that $\mathfrak{L}\subset\mathfrak{C}\backslash{\mathfrak{C}^{*}}$, which implies that $\left|J(\mathcal{C})\right|\leq s$ and that $f^{*}(\mathbf{x})=\mathbb{E}\left\{Y~|~\mathbf{x}\in\mathcal{C}\right\}$ for all $\mathcal{C}\in\mathfrak{L}$. As a consequence of (H.15), for any $\mathbf{x}\in\{\pm 1\}^{d}$ and letting $\mathcal{C}(\mathbf{x})$ be the leaf that contains it, we have
$\begin{split}\left|\hat{f}_{\operatorname{CART}}(\mathbf{x})-f^{*}(\mathbf{x})\right|&=\left|\bar{Y}_{I(\mathcal{C})}-f^{*}(\mathbf{x})\right|\\ &=\left|\bar{Y}_{I(\mathcal{C})}-\mathbb{E}\left\{Y~|~\mathbf{X}\in\mathcal{C}(\mathbf{x})\right\}\right|\\ &\leq\sqrt{\frac{2(\|f^{*}\|_{\infty}^{2}+\sigma^{2})\left((s+1)\log 3+\log(2/\delta)\right)}{N(\mathcal{C}(\mathbf{x}))}}\\ &\leq\sqrt{\frac{2^{s+1}(\|f^{*}\|_{\infty}^{2}+\sigma^{2})\left((s+1)\log 3+\log(2/\delta)\right)}{n}}.\end{split}$ (H.28)
@@
$R\left(\hat{f}_{\operatorname{CART}}(-;\mathcal{D}_{n}),f^{*}\right)\leq\frac{2^{s+1}(\|f^{*}\|_{\infty}^{2}+\sigma^{2})\left((s+1)\log 3+\log(2/\delta)\right)}{n},$ (H.29)
@@
To prove (H.25), we take $\delta=1/n$ in (H.24). Let $\mathcal{E}$ be the event on which (H.24) holds. We then write
$\mathfrak{R}\left(\hat{f}_{\operatorname{CART}},f^{*}_{0},d,n\right)=\mathbb{E}\left\{R\left(\hat{f}_{\operatorname{CART}}(-;\mathcal{D}_{n}),f^{*}\right)\mathbb{1}_{\mathcal{E}}\right\}+\mathbb{E}\left\{R\left(\hat{f}_{\operatorname{CART}}(-;\mathcal{D}_{n}),f^{*}\right)\mathbb{1}_{\mathcal{E}^{c}}\right\}.$ (H.30)
@@
$\|\hat{f}_{\operatorname{CART}}(-;\mathcal{D}_{n})\|_{\infty}\leq\max_{1\leq i\leq n}\left|Y_{i}\right|,$ (H.31)
@@
$R\left(\hat{f}_{\operatorname{CART}}(-;\mathcal{D}_{n}),f^{*}\right)\leq\left(\max_{1\leq i\leq n}\left|Y_{i}\right|+\left\|f^{*}\right\|_{\infty}\right)^{2}.$ (H.32)
@@
$\max_{1\leq i\leq n}\left|Y_{i}\right|\lesssim\sigma\left(\sqrt{\log n}+t\right)$ (H.33)
with probability at least $1-2e^{-t^{2}}$, for any $t>0$. Plugging this into (H.32) and applying Cauchy-Schwarz twice gives
$R\left(\hat{f}_{\operatorname{CART}}(-;\mathcal{D}_{n}),f^{*}\right)\lesssim\left\|f^{*}\right\|_{\infty}^{2}+\sigma^{2}\log n+\sigma^{2}u$ (H.34)
@@
$\mathbb{P}\left\{R\left(\hat{f}_{\operatorname{CART}}(-;\mathcal{D}_{n}),f^{*}\right)\geq u\right\}\leq 2e^{-C(u-\|f^{*}\|_{\infty}^{2}+\sigma^{2}\log n)/\sigma^{2}}$ (H.35)
@@
$\begin{split}\mathbb{E}\left\{R\left(\hat{f}_{\operatorname{CART}}(-;\mathcal{D}_{n}),f^{*}\right)\mathbb{1}_{\mathcal{E}^{c}}\right\}&=\int_{0}^{\infty}\mathbb{P}\left\{R\left(\hat{f}_{\operatorname{CART}}(-;\mathcal{D}_{n}),f^{*}\right)\mathbb{1}_{\mathcal{E}^{c}}\geq u\right\}du\\ &\leq\int_{0}^{\infty}\min\left\{\mathbb{P}\left\{R\left(\hat{f}_{\operatorname{CART}}(-;\mathcal{D}_{n}),f^{*}\right)\geq u\right\},\mathbb{P}\left\{\mathcal{E}^{c}\right\}\right\}du\\ &\leq\int_{0}^{\infty}\min\left\{\frac{1}{n},2e^{-C(u-\|f^{*}\|_{\infty}^{2}+\sigma^{2}\log n)/\sigma^{2}}\right\}du\\ &\lesssim\frac{\sigma^{2}\log n+\left\|f^{*}\right\|_{\infty}^{2}}{n}.\end{split}$ (H.36)
This completes the bound on the second term and hence the proof of (H.25). ∎
Appendix I Proof of Theorem 8.1
@@
Imitating the proof of Theorem 5.1, we combine Lemma D.2 together with an extension of Lemma D.1 to accommodate a vertex cut and the new coupling argument. ∎
Lemma I.1.
@@
$\operatorname{KL}\left(\mathcal{D}_{n}\|\check{\mathcal{D}}_{n}\right)=\sum_{i=1}^{n}\operatorname{KL}\left((\mathbf{X}_{i},Y_{i})\|(\mathbf{X}_{i},\check{Y}_{i})\right).$ (I.1)
@@
$\operatorname{KL}\left((\mathbf{X}_{i},Y_{i})\|(\mathbf{X}_{i},\check{Y}_{i})\right)=\operatorname{KL}(\mathbf{X}_{i}\|\mathbf{X}_{i})+\mathbb{E}\left\{\operatorname{KL}\left((Y_{i}|\mathbf{X}_{i})\|(\check{Y}_{i}|\mathbf{X}_{i})\right)\right\}.$ (I.2)
The first term on the right hand side is equal to 0 since the covariate distributions are the same. To compute the second term, notice that $Y_{i}|\mathbf{X}_{i}\sim\mathcal{N}(f^{*}(\mathbf{X}_{i}),\sigma^{2})$, and $\check{Y}_{i}|\mathbf{X}_{i}\sim\mathcal{N}(\check{f}^{*}(\mathbf{X}_{i}),\sigma^{2})$. The KL divergence between them is thus
$\operatorname{KL}\left((Y_{i}|\mathbf{X}_{i})\|(\check{Y}_{i}|\mathbf{X}_{i})\right)=\frac{\left(f^{*}(\mathbf{X}_{i})-\check{f}^{*}(\mathbf{X}_{i})\right)^{2}}{2\sigma^{2}}.$ (I.3)
@@
$\mathbb{E}\left\{\operatorname{KL}\left((Y_{i}|\mathbf{X}_{i})\|(\check{Y}_{i}|\mathbf{X}_{i})\right)\right\}=\frac{\mathbb{E}\left\{\left(\sum_{S\in\mathcal{B}}\alpha_{S}\chi_{S}(\mathbf{X}_{i})\right)\right\}^{2}}{2\sigma^{2}}=\frac{w(\mathcal{B})}{2\sigma^{2}}.$ (I.4)
Plugging this back into (I.1) gives
$\operatorname{KL}\left(\mathcal{D}_{n}\|\check{\mathcal{D}}_{n}\right)\leq\frac{nw(\mathcal{B})}{2\sigma^{2}},$ (I.5)
@@
$\operatorname{TV}\left(\mathcal{D}_{n},\check{\mathcal{D}}_{n}\right)\leq\sqrt{\frac{nw(\mathcal{B})}{4\sigma^{2}}}.$ (I.6)
@@
$\begin{split}\mathbb{P}_{\mathcal{D}_{n}}\left\{J(\mathbf{x};\mathcal{D}_{n},\Theta)\cap T\neq\emptyset\right\}&\leq\mathbb{P}_{\mathcal{D}_{n}}\left\{J(\mathbf{x};\check{\mathcal{D}}_{n},\Theta)\cap T\neq\emptyset~\text{or}~\check{\mathcal{D}}_{n}\neq\mathcal{D}_{n}\right\}\\ &\leq\mathbb{P}_{\mathcal{D}_{n}}\left\{J(\mathbf{x};\check{\mathcal{D}}_{n},\Theta)\cap T\neq\emptyset\right\}+\mathbb{P}\left\{\check{\mathcal{D}}_{n}\neq\mathcal{D}_{n}\right\}\\ &\frac{s_{\operatorname{T}}(\log_{2}n+2)}{d-s_{-\mathcal{B},\textsf{MSP}}-s_{\operatorname{T}}+1}+\sqrt{\frac{nw(\mathcal{B})}{4\sigma^{2}}}.\qed\end{split}$ (I.7)
Appendix J Constant Probability Lower Bounds
@@
$R\left(\hat{f}_{\operatorname{G}}(-;\mathcal{D}_{n},\Theta),f^{*}\right)\geq\frac{\operatorname{Var}\left\{r_{\textsf{MSP}}(\mathbf{X})\right\}}{6}$ (J.1)
@@
$\log_{2}n\leq\frac{\operatorname{Var}\{r_{\textsf{MSP}}(\mathbf{X})\}(d-s_{\textsf{MSP}}-s_{\operatorname{T}}+1)}{8M^{2}s_{\textsf{MSP}}}-2,$ (J.2)
@@
$\mathbb{E}_{\Theta}\left\{R\left(\hat{f}_{\operatorname{G-E}}(-;\mathcal{D}_{n},\Theta)\right)\right\}\geq\frac{\operatorname{Var}\left\{r_{\textsf{MSP}}(\mathbf{x})\right\}}{4}.$ (J.3)
@@
$R\left(\hat{f}_{\operatorname{G}}(-;\mathcal{D}_{n},\Theta),f^{*}\right)\geq V(\mathcal{D}_{n},\Theta)\operatorname{Var}\left\{r_{\textsf{MSP}}(\mathbf{X})\right\},$ (J.4)
@@
[…]
It is clear that $0\leq V(\mathcal{D}_{n},\Theta)\leq 1$. Furthermore, using Lemma D.1, we see that $\mathbb{E}_{\mathcal{D}_{n}}\left\{V(\mathcal{D}_{n},\Theta)\right\}\geq 1-\delta$, where $\delta\coloneqq\frac{s_{\textsf{MSP}}(\log_{2}n+2)}{d-s_{\textsf{MSP}}-s_{\operatorname{T}}+1}$. Next, making use of the simple inequality
$\mathbb{E}_{\mathcal{D}_{n}}\left\{V(\mathcal{D}_{n},\Theta)\right\}\leq\mathbb{P}_{\mathcal{D}_{n}}\left\{V(\mathcal{D}_{n},\Theta)\geq\delta\right\}+\delta,$ (J.6)
@@
$\mathbb{P}_{\mathcal{D}_{n}}\left\{V(\mathcal{D}_{n})\geq\delta\right\}\geq 1-3\delta.$ (J.7)
@@
$R\left(\hat{f}_{\operatorname{G}}(-;\mathcal{D}_{n},\Theta),f^{*}\right)\geq\frac{\operatorname{Var}\left\{r_{\textsf{MSP}}(\mathbf{X})\right\}}{6}.$ (J.8)